# BD-14 3065b (TOI-4987b): from giant planet to brown dwarf: evidence for deuterium burning in old age? Key Words.: giant planets – brown dwarfs – techniques: photometric – techniques: spectroscopic – techniques: radial velocities1 Introduction

### v1

@@
The present study reports the confirmation of BD-14 3065b, a transiting planet/brown dwarf in a triple-star system, with a mass near the deuterium burning boundary. BD-14 3065b has the largest radius observed within the sample of giant planets and brown dwarfs around post-main-sequence stars. Its orbital period is 4.3 days, and it transits a subgiant F-type star with a mass of M, a radius of R, an effective temperature of K, and a metallicity of dex. By combining TESS photometry with high-resolution spectra acquired with the TRES and Pucheros+ spectrographs, we measured a mass of and a radius of . Our discussion of potential processes that could be responsible for the inflated radius led us to conclude that deuterium burning is a plausible explanation resulting from the heating of BD-14 3065b’s interior. Detection of the secondary eclipse with TESS photometry enables a precise determination of the eccentricity and reveals BD-14 3065b has a brightness temperature of K. With its unique characteristics, BD-14 3065b presents an excellent opportunity to study its atmosphere through thermal emission spectroscopy.
@@ 2.1 TESS light curves
Green triangles indicate the positions of transits. The LCs still show pulsations, but these were later removed.
@@ 2.2 Contamination from nearby sources
As part of follow-up observations coordinated by the TESS Follow-up Observing Program (TFOP) High-Resolution Imaging Sub-Group 3 (SG3), we obtained high angular-resolution imaging of BD-14 3065. Observations were made on April 15, 2022, using the High-Resolution Camera (HRCam; Tokovinin & Cantarutti, 2008) speckle interferometry instrument on the Southern Astrophysical Research (SOAR) 4.1m telescope. The observation strategy and data reduction procedures are described in Tokovinin (2018), or Ziegler et al. (2021). Observations were made in the Cousins filter with a resolution of 36 mas. The final reconstructed image, shown in Fig. 4, reaches a contrast of at a separation of 0.5 arcsec and has an estimated PSF that is 0.064 arcsec wide. The speckle imaging of BD-14 3065 detected a companion at an angular separation of 0.92 arcsec from the primary (translated into a projected physical separation of 520 au), which has mag, and position angle of 210.5 or 30.5 degrees. We note that Gaia did not report this companion. However, the Gaia Renormalised Unit Weight Error (RUWE) value of 3.5 for BD-14 3065 suggests the presence of a close multiplicity. We discuss the treatment of the companion in Section 3.1.
@@ 2.3 Ground-based light curves
As part of the TFOP, ground-based photometric data were collected for BD-14 3065. The observations were scheduled using Transit Finder, a customized version of the Tapir software (Jensen, 2013), and the photometric data was extracted using AstroImageJ (Collins et al., 2017). However, the light curve precision is not sufficient to improve the parameters from the TESS LCs. Moreover, with some precision, different dilution parameters would need to be adopted as these observations were made in different filters. Therefore, it was decided not to consider ground-based photometry in this paper further.
RV (m s) | (m s) | Instrument | |
2459601.934527 | 5017 | 144 | TRES |
2459605.955459 | 4746 | 111 | TRES |
2459616.890452 | 7113 | 144 | TRES |
2459617.956610 | 5910 | 162 | TRES |
2459618.882581 | 4777 | 155 | TRES |
2459619.825369 | 4944 | 101 | TRES |
2459620.899334 | 6977 | 143 | TRES |
2459621.881599 | 6721 | 116 | TRES |
2459622.928761 | 4436 | 87 | TRES |
2459623.911069 | 4917 | 112 | TRES |
2459624.901062 | 6643 | 105 | TRES |
2459625.944000 | 6830 | 125 | TRES |
2459626.800265 | 5017 | 107 | TRES |
2459635.850469 | 4864 | 168 | TRES |
2459922.954247 | 1358 | 172 | TRES |
2459923.986662 | 257 | 173 | TRES |
2459924.995943 | 2389 | 156 | TRES |
2459930.979474 | 2002 | 113 | TRES |
2459932.988929 | 252 | 93 | TRES |
2459933.984636 | 2123 | 141 | TRES |
2459934.972431 | 2169 | 156 | TRES |
2459936.000038 | 32 | 102 | TRES |
2459936.965439 | -380 | 122 | TRES |
2459944.924846 | 37 | 160 | TRES |
2459948.961852 | 67 | 107 | TRES |
2459957.965266 | -70 | 152 | TRES |
2459963.949222 | 2463 | 205 | TRES |
2460013.776225 | 57 | 153 | TRES |
2460014.767013 | 1006 | 119 | TRES |
2460016.749179 | 1564 | 113 | TRES |
2460017.788237 | -82 | 94 | TRES |
2460028.772600 | 1471 | 129 | TRES |
2460029.856313 | 1056 | 127 | TRES |
2460030.758805 | 32 | 145 | TRES |
2460071.660621 | 1984 | 196 | TRES |
2460072.653619 | 1507 | 133 | TRES |
2460073.691053 | 457 | 129 | TRES |
2460074.681585 | 962 | 140 | TRES |
2460075.670057 | 2148 | 118 | TRES |
2460076.660340 | 2017 | 142 | TRES |
2460077.697400 | 721 | 198 | TRES |
2460077.486053 | 2324 | 1122 | Pucheros+ |
2460078.557373 | 715 | 1114 | Pucheros+ |
2460078.580683 | 2164 | 1455 | Pucheros+ |
2460080.569803 | 3522 | 973 | Pucheros+ |
2460080.590706 | 3201 | 1857 | Pucheros+ |
2460085.523472 | 3111 | 867 | Pucheros+ |
2460085.544838 | 4610 | 901 | Pucheros+ |
2460086.569549 | 1530 | 1006 | Pucheros+ |
2460086.590926 | 2332 | 1117 | Pucheros+ |
2460088.507060 | 5548 | 1197 | Pucheros+ |
2460088.528426 | 3275 | 1204 | Pucheros+ |
2460089.519780 | 4380 | 1107 | Pucheros+ |
2460094.491713 | 3257 | 794 | Pucheros+ |
2460096.531829 | 3333 | 374 | Pucheros+ |
2460102.509028 | 4203 | 551 | Pucheros+ |
@@ 2.4 TRES spectra
We obtained spectra of BD-14 3065 using the Tillinghast Reflector Echelle Spectrograph (TRES) on the 1.5-m telescope at the Whipple Observatory on Mount Hopkins, Arizona, from January 22, 2022 to May 13, 2023. The spectrograph has a resolving power of R 44 000 and covers wavelengths ranging from 390 nm to 910 nm. A total of 41 spectra of BD-14 3065 were captured with TRES, with exposure times varying from 350–3600 s depending on the weather conditions. The S/N ratio per resolution element of these spectra ranged from 24 to 49. We derived the relative RVs from the TRES spectra by considering multiple echelle orders from each spectrum, which were cross-correlated with a high SNR median spectrum produced using the complete set of observed spectra. We have also adjusted the spectra to account for the small drifts in the TRES zero point over the period of the observations. The TRES fiber diameter measures 2.4 arcseconds and contamination from a stellar companion around BD-14 3065, as resolved by speckle imaging, may be expected. The exact level of contamination also depends on the seeing. The final list of radial velocities (RVs) can be found in Table 2.
@@ 2.5 Spectroscopy from the PLATOSpec project
We acquired additional spectra from the PLATOSpec project using the fully operational telescope E152 at La Silla, Chile. An interim spectrograph, PUCHEROS+, has been installed for scientific use. PUCHEROS+ is an improved version of a PUCHEROS spectrograph (Vanzi et al., 2012), with a resolving power of R . The Andor iKon M detector provides a wavelength range of over 400 to 700 nm. Typically, the spectrograph’s radial velocity stability is about 100 m s over a month for cool main sequence stars with V mag. PUCHEROS+ uses the ceres+ pipeline (Brahm et al., 2017), which produces wavelength calibrated and optimally extracted order-by-order spectra. The radial velocities are obtained using the cross-correlation method from the pipeline. The PUCHEROS+ fiber diameter has 2.0 arcseconds, and the expected contamination from a stellar companion around BD-14 3065 is smaller than for the TRES spectra. While the data will be available in the ESO archive later, they are currently accessible upon request from the PLATOSpec team. The target was observed over the course of 16 nights between May and June 2023, with each exposure lasting 1800 seconds.
@@ 3.1 Modeling Stellar Parameters
Parameter BD-14 3065 A BD-14 3065 B BD-14 3065 AB () - () - - (K) - - (km s) - - (days) - - log L () - log Age (Gyr) - (mag) - Galex mag Galex mag Tycho-2 mag Gaia mag Tycho-2 mag Gaia mag Gaia mag TESS mag 2MASS mag 2MASS mag 2MASS mag WISE1 mag WISE2 mag WISE3 mag Parameter BD-14 3065 A BD-14 3065c BD-14 3065 B () () (K) (km s) - - (days) - - log L () log Age (Gyr) (mag) Galex mag Galex mag Tycho-2 mag Gaia mag Tycho-2 mag Gaia mag Gaia mag TESS mag 2MASS mag 2MASS mag 2MASS mag WISE1 mag WISE2 mag WISE3 mag
@@
The broad-band catalog and time-series photometry of BD-14 3065 need to be corrected for the contamination of the second star found in the speckle imaging. Hence, we used the python package called isochrones to create a two-component model of the SED, taking into account the priors set on the derived spectral parameters, available catalog photometry, mag difference from speckle imaging, and Gaia parallax. We also considered the line-of-sight extinction by placing an upper limit derived from the Bayestar dust maps of Green et al. (2019) incorporated into the dustmaps code (Green, 2018). We have obtained the stellar parameters of both stars reported in Table 3.
@@
We performed a test to determine if the G-type companion star, discovered through speckle imaging, impacts our spectral parameters. To do this, we synthesized the spectrum of a G-type star with a of 2.3 mag. In the wavelength range of nm, the mag is 2.5, indicating that the companion star is ten times fainter. However, since a G-type star has deeper spectral lines compared to the F-type primary star, the companion star can still affect our findings. We also considered that a G-type star has a slower rotation rate. Figure 5 shows the spectra of both stars. Our findings revealed that the contamination could cause an underestimation of the effective temperature by 200 K and an overestimation of the value of by 0.1 dex. It is consistent with the work of El-Badry et al. (2018), who studied the implications of unresolved binaries in stellar spectra for spectral fitting. They presented a figure that describes the effects of unresolved binarity on the accuracy of stellar labels recovered from spectral fitting as a function of the mass ratio between primary and secondary. Therefore, we adjusted our priors placed on the derived spectral parameters in SED modeling to take this effect into account. We plot the best-fit SED model in the Fig. 6, and present the stellar properties of the primary and secondary components in Table 3. We also overplot in Fig. 7 luminosities and effective temperatures of both stars with the MIST stellar evolutionary tracks (Choi et al., 2016). According to the analysis of the stellar isochrones, the system appears to have an age of approximately Gyr. In order to corroborate this finding, we searched for the lithium nm feature in the co-added TRES spectrum, yet did not detect any indication of it, further supporting the system’s old age. Additionally, the synchronized rotation period of the star, found in Section 3.1.2, with the planet’s orbital period suggests tidal evolution, which also aligns with the old age of the system. Finally, the PyAstronomy python package was utilized to determine the ,, velocities for BD-14 3065, using the gal_uvw222https://pyastronomy.readthedocs.io/en/latest/pyaslDoc/aslDoc/gal_uvw.html function. Based on our analysis, we found that BD-14 3065 is associated with the thin disk (e.g., Leggett, 1992; Bensby et al., 2003). This outcome is predictable since the thick disk is mainly composed of stars that are at least 8 Gyr old (e.g., Fuhrmann, 1998).
@@ 3.1.2 Stellar rotation
where variables and define the amplitude of the GP, while represents the time-lag and denotes the amplitude-modulation timescale of the GP. indicates the rotational period. The maximum a posteriori (MAP) parameters are estimated using the L-BFGS-B nonlinear optimization routine (Byrd et al., 1995; Zhu et al., 1997). To obtain the marginalized posterior distributions of free parameters, we run MCMC using emcee (Goodman & Weare, 2010; Foreman-Mackey et al., 2013). The analysis results in a derived rotational period of days, which is shown in Fig. 10 as the probability density of alongside the MAP model prediction. Additionally, we apply the generalized Lomb-Scargle (GLS) periodograms (Zechmeister & Kürster, 2009) to the TESS LCs as an independent check of the derived . We observe strong peaks around 4.3 days in individual sectors and a forest of peaks around 4.3 days when using all sectors together. We consider this to be the rotation period, and the periodogram is shown in Fig. 10. Table 3 reports the derived rotation period. By using combination of , , and we can constrain stellar inclination angle, which is for the two-star model and for the three-star model. Our approach follows the procedure described in Masuda & Winn (2020) and used in Bowler et al. (2023). This method properly accounts for the correlation between stellar equatorial velocity and projected rotational velocity. Posterior distributions of stellar inclination for both models can be seen in Fig. 11. It should be noted that the used value of represents an upper limit, as the spectral contamination of BD-14 1065 can lead to an overestimation of the value by a few km s. Correction for this overestimation would lead to an even higher probability of spin-orbit misalignment.
@@ 3.2 System architecture
The high brightness ratio of two stars posed significant challenges to revealing the true nature of the BD-14 3065 system. To do so, we adopted the approach detailed in Mandushev et al. (2005). Initially, we computed the bisector spans (BIS) of cross-correlation functions (CCFs) to examine the asymmetries of spectral lines. We then determined the periodogram of the BIS time series (see Section 3.3), which revealed two peaks at 4.3 and 3.3 days. These peaks indicated that two CCFs are moving with respect to each other with the period of orbital motion, causing asymmetries in the blended CCF. However, here, we observed a key difference between the GSC01944-02289 and BD-14 3065 systems. While BIS showed a semiamplitude of more than one in the case of GSC01944-02289, BD-14 3065 BIS only showed a semiamplitude of less than 300 , and the peak was barely significant, albeit visible.
@@
Subsequently, we performed the two-dimensional cross-correlation algorithm TODCOR (Zucker & Mazeh, 1994). TODCOR cross-correlates the observed spectra against a composite template obtained by adding together two distinct templates chosen to match each star. We used stellar parameters from the two-star SED modeling and assumed that the stars in the potential eclipsing binary would likely have their rotation synchronized with the orbital motion due to tidal forces (translating to the vsini of 11 ). We were able to measure RVs of both stars. An illustration of this detection is given in Fig. 12, which shows the two-dimensional correlation function as well as cross-sections of the two-dimensional correlation function at a fixed velocity for the G star and F star, respectively. The periodogram of the F-type star (see Section 3.3) confirmed the substellar nature of the transiting object.
@@
To further validate our findings, we conducted additional tests. Firstly, we allowed the light fraction factor to be a free parameter, and the results indicated that the best value was close to the expected value of 0.1 from speckle imaging. Secondly, we recomputed the periodogram of BIS from the cross-section of 2D CCF of the F-type star. This revealed a significant reduction in the scatter of data, which was now on the same level as observed in the Sun. It also eliminated the peak at the period of orbital motion, suggesting that we had successfully extracted the contamination of the G-type star. Finally, we compared the photometry of GSC01944-02289 and BD-14 3065. In the TESS LCs of GSC01944-02289, clear variations with the period of half the orbital period of the eclipsing star are visible, with maxima in phases 0.25 and 0.75. The stellar deformation of close binary stars causes these variations. We would, therefore, expect to observe similar variations for BD-14 3065. However, the LCs instead show variations with a period similar to the transiting companion’s period and a clear evolution of this variability through TESS sectors, which we interpret as the rotation period of the F-type star (see Fig. 1). Such a rotation period would be an improbable coincidence if the companion orbited a different star and no tidal interactions were in play. In conclusion, the transiting object transits an F-type primary star rather than a wide stellar companion.

### v2

@@
The present study reports the confirmation of BD-14 3065b, a transiting planet/brown dwarf in a triple-star system, with a mass near the deuterium burning boundary. BD-14 3065b has the largest radius observed within the sample of giant planets and brown dwarfs around post-main-sequence stars. Its orbital period is 4.3 days, and it transits a subgiant F-type star with a mass of M⊙, a radius of R⊙, an effective temperature of K, and a metallicity of dex. By combining TESS photometry with high-resolution spectra acquired with the TRES and Pucheros+ spectrographs, we measured a mass of and a radius of . Our discussion of potential processes that could be responsible for the inflated radius led us to conclude that deuterium burning is a plausible explanation resulting from the heating of BD-14 3065b’s interior. Detection of the secondary eclipse with TESS photometry enables a precise determination of the eccentricity and reveals BD-14 3065b has a brightness temperature of K. With its unique characteristics, BD-14 3065b presents an excellent opportunity to study its atmosphere through thermal emission spectroscopy.
@@ 2.2 Contamination from nearby sources
As part of follow-up observations coordinated by the TESS Follow-up Observing Program (TFOP) High-Resolution Imaging Sub-Group 3 (SG3), we obtained high angular-resolution imaging of BD-14 3065. Observations were made on April 15, 2022, using the High-Resolution Camera (HRCam; Tokovinin & Cantarutti, 2008) speckle interferometry instrument on the Southern Astrophysical Research (SOAR) 4.1m telescope. The observation strategy and data reduction procedures are described in Tokovinin (2018), or Ziegler et al. (2021). Observations were made in the Cousins filter with a resolution of 36 mas. The final reconstructed image, shown in Fig. 4, reaches a contrast of at a separation of 0.5 arcsec and has an estimated PSF that is 0.064 arcsec wide. The speckle imaging of BD-14 3065 detected a companion at an angular separation of 0.92 arcsec from the primary (translated into a projected physical separation of 520 au), which has mag, and position angle of 210.5 or 30.5 degrees. In order to conduct further analyses, we adopted a conservative errorbar of 0.1 mag for magnitude difference. We also note that Gaia did not report this companion. However, the Gaia Renormalised Unit Weight Error (RUWE) value of 3.5 for BD-14 3065 suggests the presence of a close multiplicity. We discuss the treatment of the companion in Section 3.1.
@@ 2.3 Ground-based light curves
As part of the TFOP, ground-based photometric data were collected for BD-14 3065. The observations were scheduled using Transit Finder, a customized version of the Tapir software (Jensen, 2013), and the photometric data was extracted using AstroImageJ (Collins et al., 2017).
Our team conducted four transits using the Las Cumbres Observatory Global Telescope (LCOGT; Brown et al., 2013) 1.0-m network. The telescopes are equipped with 4096×4096 pixel SINISTRO cameras that have an image scale of 0.389 arcsec per pixel. Two transits were observed in alternating filter mode, resulting in a total of four light curves. Together, three transits are in the Sloan filter, and three transits in the Sloan filter. The images were calibrated using the standard LCOGT BANZAI pipeline (McCully et al., 2018). The aperture radius used ranges from 5.5 to 10.5 arcsec.
Additionally, one transit was observed with the 0.305m Perth Exoplanet Survey Telescope (PEST) in the Rc filter. The telescope is equipped with a QHY183M camera with an image scale of 0.71 arcsec per pixel. The data reduction and aperture photometry were performed using a custom software package PEST Pipeline222https://pestobservatory.com/the-pest-pipeline/. The aperture radius used for the observations was 5.7 arcsec.
@@ 2.4 TRES spectra
We obtained spectra of BD-14 3065 using the Tillinghast Reflector Echelle Spectrograph (TRES) on the 1.5-m telescope at the Whipple Observatory on Mount Hopkins, Arizona, from January 22, 2022 to April 13, 2024. The spectrograph has a resolving power of R 44 000 and covers wavelengths ranging from 390 nm to 910 nm. A total of 57 spectra of BD-14 3065 were captured with TRES, with exposure times varying from 300–3600 s depending on the weather conditions. The S/N ratio per resolution element of these spectra ranged from 24 to 60. We derived the relative RVs from the TRES spectra by considering multiple echelle orders from each spectrum, which were cross-correlated with a high SNR median spectrum produced using the complete set of observed spectra. We have also adjusted the spectra to account for the small drifts in the TRES zero point over the period of the observations. The TRES fiber diameter measures 2.4 arcseconds, and contamination from the wide stellar companion around BD-14 3065A, as resolved by speckle imaging, may be expected. The exact level of contamination also depends on the seeing. The final list of radial velocities (RVs) can be found in Table LABEL:tab:long.
@@ 2.5 Spectroscopy from the PLATOSpec project
We acquired additional spectra from the PLATOSpec project using the fully operational telescope E152 at La Silla, Chile. An interim spectrograph, PUCHEROS+, has been installed for scientific use. PUCHEROS+ is an improved version of a PUCHEROS spectrograph (Vanzi et al., 2012), with a resolving power of R . The Andor iKon M detector provides a wavelength range of over 400 to 700 nm. Typically, the spectrograph’s radial velocity stability is about 100 m s-1 over a month for cool main sequence stars with V mag. PUCHEROS+ uses the ceres+ pipeline (Brahm et al., 2017), which produces wavelength calibrated and optimally extracted order-by-order spectra. The radial velocities are obtained using the cross-correlation method from the pipeline. The PUCHEROS+ fiber diameter has 2.0 arcseconds, and expected contamination from the wide stellar companion around BD-14 3065A is much smaller than for the TRES spectra. While the data will be available in the ESO archive later, they are currently accessible upon request from the PLATOSpec team. A total of 34 spectra of BD-14 3065 were captured between May 2023 and March 2024, with each exposure lasting 1800 seconds.
@@ 3.1 Modeling Stellar Parameters
Parameter BD-14 3065 A BD-14 3065 B BD-14 3065 AB () - () - - (K) - - (km s-1) - - (days) - - log L () - log Age (Gyr) - (mag) - Galex mag Galex mag Tycho-2 mag Gaia mag Tycho-2 mag Gaia mag Gaia mag TESS mag 2MASS mag 2MASS mag 2MASS mag WISE1 mag WISE2 mag WISE3 mag Parameter BD-14 3065 A BD-14 3065c BD-14 3065 B () () (K) (km s-1) - - (days) - - log L () log Age (Gyr) (mag) Galex mag Galex mag Tycho-2 mag Gaia mag Tycho-2 mag Gaia mag Gaia mag TESS mag 2MASS mag 2MASS mag 2MASS mag WISE1 mag WISE2 mag WISE3 mag
@@
The broad-band catalog and time-series photometry of BD-14 3065 need to be corrected for the contamination of the second star found in the speckle imaging. Hence, we used the python package called isochrones to create a two-component model of the SED, taking into account the priors set on the derived spectral parameters, available catalog photometry, mag difference from speckle imaging, and Gaia parallax. We also considered the line-of-sight extinction by placing an upper limit derived from the Bayestar dust maps of Green et al. (2019) incorporated into the dustmaps code (Green, 2018). We have obtained the stellar parameters of both stars reported in Table 2. The spectral types used to differentiate the stars throughout the text are based on the empirical spectral type-color sequence from Pecaut & Mamajek (2013).
@@
We performed a test to determine if the G-type companion star, discovered through speckle imaging, impacts our spectral parameters. To do this, we synthesized the spectrum of a G-type star with a of 2.3 mag. In the wavelength range of nm, the mag is 2.5, indicating that the companion star is ten times fainter. However, since a G-type star has deeper spectral lines compared to the F-type primary star, the companion star can still affect our findings. We also considered that a G-type star has a slower rotation rate. Figure 5 shows the spectra of both stars. Our findings revealed that the contamination could cause an underestimation of the effective temperature by 200 K and an overestimation of the value of by 0.1 dex. It is consistent with the work of El-Badry et al. (2018), who studied the implications of unresolved binaries in stellar spectra for spectral fitting. They presented a figure that describes the effects of unresolved binarity on the accuracy of stellar labels recovered from spectral fitting as a function of the mass ratio between primary and secondary. Therefore, we adjusted our priors placed on the derived spectral parameters in SED modeling to take this effect into account. We plot the best-fit SED model in the Fig. 6, and present the stellar properties of the primary and secondary components in Table 2. We also overplot in Fig. 7 luminosities and effective temperatures of both stars with the MIST stellar evolutionary tracks (Choi et al., 2016). According to the analysis of the stellar isochrones, the system appears to have an age of approximately Gyr. In order to corroborate this finding, we searched for the lithium nm feature in the co-added TRES spectrum, yet did not detect any indication of it, further supporting the system’s old age. Additionally, the synchronized rotation period of the star, found in Section 3.1.2, with the planet’s orbital period suggests tidal evolution, which also aligns with the old age of the system. Finally, the PyAstronomy python package was utilized to determine the ,, velocities for BD-14 3065, using the gal_uvw333https://pyastronomy.readthedocs.io/en/latest/pyaslDoc/aslDoc/gal_uvw.html function. Based on our analysis, we found that BD-14 3065 is associated with the thin disk (e.g., Leggett, 1992; Bensby et al., 2003). This outcome is predictable since the thick disk is mainly composed of stars that are at least 8 Gyr old (e.g., Fuhrmann, 1998).
@@ 3.1.2 Stellar rotation
where variables and define the amplitude of the GP, while represents the time-lag and denotes the amplitude-modulation timescale of the GP. indicates the rotational period. The maximum a posteriori (MAP) parameters are estimated using the L-BFGS-B nonlinear optimization routine (Byrd et al., 1995; Zhu et al., 1997). To obtain the marginalized posterior distributions of free parameters, we run MCMC using emcee (Goodman & Weare, 2010; Foreman-Mackey et al., 2013). The analysis results in a derived rotational period of days, which is shown in Fig. 10 as the probability density of alongside the MAP model prediction. Additionally, we apply the generalized Lomb-Scargle (GLS) periodograms (Zechmeister & Kürster, 2009) to the TESS LCs as an independent check of the derived . We observe strong peaks around 4.3 days in individual sectors and a forest of peaks around 4.3 days when using all sectors together. We consider this to be the rotation period, and the periodogram is shown in Fig. 10. Table 2 reports the derived rotation period. By using combination of , , and we can constrain stellar inclination angle, which is degrees for the two-star model and degrees for the three-star model. Our approach follows the procedure described in Masuda & Winn (2020) and used in Bowler et al. (2023). This method properly accounts for the correlation between stellar equatorial velocity and projected rotational velocity. Posterior distributions of stellar inclination for both models can be seen in Fig. 11. It should be noted that the used value of represents an upper limit, as the spectral contamination of BD-14 1065 can lead to an overestimation of the value by a few km s-1. Correction for this overestimation would lead to an even higher probability of spin-orbit misalignment.
@@ 3.2 System architecture
The high brightness ratio of two stars posed significant challenges to revealing the true nature of the BD-14 3065 system. To do so, we adopted the approach detailed in Mandushev et al. (2005) on the TRES spectra. Initially, we computed the bisector spans (BIS) of cross-correlation functions (CCFs) to examine the asymmetries of spectral lines. We then determined the periodogram of the BIS time series (see Section 3.3), which revealed two peaks at 4.3 and 3.3 days. These peaks indicated that two CCFs are moving with respect to each other with the period of orbital motion, causing asymmetries in the blended CCF. However, here, we observed a key difference between the GSC01944-02289 and BD-14 3065 systems. While BIS showed a semiamplitude of more than one in the case of GSC01944-02289, BD-14 3065 BIS only showed a semiamplitude of less than 300 , and the peak was barely significant, albeit visible.
@@
Subsequently, we performed the two-dimensional cross-correlation algorithm TODCOR (Zucker & Mazeh, 1994). TODCOR cross-correlates the observed spectra against a composite template obtained by adding together two distinct templates chosen to match each star. We used stellar parameters from the two-star SED modeling and assumed that the stars in the potential eclipsing binary would likely have their rotation synchronized with the orbital motion due to tidal forces (translating to the vsini of 11 ). We were able to measure the RVs of both stars. An illustration of this detection is given in Fig. 12, which shows the two-dimensional correlation function as well as cross-sections of the two-dimensional correlation function at a fixed velocity for the G star and F star, respectively.
Notably, the detection of the transiting object around the F star (see Figure 13) confirms its substellar nature.
@@
To further validate the results from the TODCOR analysis, we conducted additional tests. Firstly, we allowed the light fraction factor to be a free parameter, and the results indicated that the best value was close to the expected value of 0.1 from speckle imaging. Secondly, we analyzed data of BIS from the cross-section of 2D CCF of the F-type star and computed its periodogram. This revealed that the magnitude of the asymmetries in the BIS data decreased. It also eliminated the peak at the period of orbital motion, suggesting that we had successfully extracted the contamination of the G-type star. Importantly, the Pucheros+ data are less affected by the wide stellar companion than TRES data. Therefore, the presence of the signal of the transiting planet in this dataset indicates that it originates in the primary star. Finally, we compared the photometry of GSC01944-02289 and BD-14 3065. In the TESS LCs of GSC01944-02289, clear variations with the period of half the orbital period of the eclipsing star are visible, with maxima in phases 0.25 and 0.75. The stellar deformation of close binary stars causes these variations. We would, therefore, expect to observe similar variations for BD-14 3065. However, the LCs instead show variations with a period similar to the transiting companion’s period and a clear evolution of this variability through TESS sectors, which we interpret as the rotation period of the F-type star (see Fig. 1). It is noteworthy that most of the flux in the TESS photometry is coming from the F star, and if these photometric variations originated from the G star, they would be suppressed by the flux from the F star. It would be surprising for an old G-type star to exhibit such large fluctuations. If the companion orbited the G star and no tidal interactions were in play, the close rotation period of the F star would be an unlikely coincidence.
In conclusion, using the TODCOR tool, we were able to distinguish F and G star components in the cross-correlation function. Our additional validations revealed that the analysis was successful and the transiting object is around the F star. Moreover, the signal of the transiting planet is also visible in less contaminated Pucheros+ data and our discussion suggests that the TESS photometry aligns with this conclusion.
As an alternative approach, ground-based photometry may be utilized to differentiate which star the transiting body orbits. Our calculations indicate that if the body orbits the primary F star, transit depths in the Sloan filter are expected to be only 0.1 ppt deeper than in the Sloan filter. However, such a small difference would not be distinguishable in the LCOGT photometry. In the event that the body transits one of a fainter star, transit depths in the Sloan filter would be roughly 1 ppt deeper than in the Sloan filter. With applied up to two detrending vectors to the LCOGT photometry, we obtained a good solution for both scenarios, one in which transit depths are similar in both filters and another with transit depth shallower in the Sloan filter. Due to the limited number of out-of-transit observations, we were unable to investigate the impact of systematics on the transit depth. Nevertheless, by computing the Bayesian model log evidence (ln Z), a widely used metric, we can distinguish the scenario with the same depth transits as significantly better. In Figure A.22 we plot the scenario with all transists having a similar depth.